OpenStudy (anonymous):

ltwelve-seventeenths divided by four-seventeenths forty-eight seventeen 4 3

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

Is that multiple choice?

OpenStudy (anonymous):

yes but its has to be correct

OpenStudy (anonymous):

= 12/17 divided by 4/17 You can just cross multiply; = 12/17 x 17/4 = 204/68 = 3

OpenStudy (anonymous):

12/17 X 17/4

OpenStudy (anonymous):

you get it..?

OpenStudy (anonymous):

How can he not get it when you gave him the answer?

OpenStudy (anonymous):

i am a she not a he and yes i get it

OpenStudy (anonymous):

medal plz..??

OpenStudy (anonymous):

and how she get it @openstudy14 when you are stucking his mind in unusual way.????

OpenStudy (anonymous):

@zaibali.qasmi I was just trying to lead her up to how to solve it without giving her the answer.